# Partial persistence of memory in bubble breakup: incomplete universality acquired by broken symmetry

### v8

@@ Abstract
When a water drop falls from a faucet, the drop is created with the formation of an axisymmetric constriction region, which thins down to breakup. Such formation of a fluid drop has been extensively studied as a representative case of the singular dynamics widely observed in nature. The singular dynamics is often self-similar: shape at different times collapsing onto a master curve after rescaling. The self-similar dynamics has been categorized as either universal or non-universal: the master curve is independent from or dependent on the length scales that set the initial boundary conditions, as if memory is erased or retained. In the previously known cases, where the axisymmetry is maintained, only a single length scale is available for the boundary conditions, a tube radius, and, thus, there is no possibility of partial memory. Here, we focus on the post-breakup and confine the system to break the axisymmetry, introducing three length scales, which leads to a third category of incomplete universality, where memory is partially retained: the master curve for the post-breakup is dependent on the smallest scale but independent from the other two scales. Affecting of only the smallest length scale on the master curve underscores the importance of scale separation for the emergence of universality: near the singularity physics at small scales becomes important. The present study suggests a promising direction for the study on the singular dynamics by exploring the symmetry and has implications for bubble/drop generation in confined geometries, encountered in various fields from microfluidics to geology.
Singular dynamics observed widely in nature from the gravitational collapse of a star [1] to the formation of a drop in dripping faucets [2] have attracted scientist in many fields such as cosmology [3], hydrodynamics, physics and mathematics [4, 5]. One important feature of the singular dynamics governed by partial differential equations is self-similarity or scale-invariance, i.e., the shapes at different times can be superimposed onto a master curve by rescaling, as observed in various phenomena including fluid-jet formation [6, 7] and drop coalescence [8, 9, 10]. The self-similar dynamics is said to be universal and losing memory, when the master curve representing the self-similarity is independent of the scales that set the initial boundary conditions with remaining dependence on only a few material parameters. This scenario is expected from separation of scales: the physics at small scales in length and time becomes so important that the dynamics does not exhibit any dependence on larger scales such as those setting the boundary conditions. One example of the self-similar and universal dynamics is the breakup of a drop surrounded by less viscous fluid, where capillarity, inertia, and viscosity all come into play [11]. However, the breakup of bubble surrounded by more viscous fluid was shown to be self-similar but non-universal, where the dynamics governed by the Stokes equation describing the viscous-capillary balance, which does not define internal length scales [12]: the master curve representing the self-similarity does depend on the boundary condition, retaining memory. In addition, the crossover from universal to non-universal self-similar dynamics is reported for the pinch-off of a bubble in a tube [13], where both regimes are governed by capillarity and viscosity and the universal regime is described by a wetting rim dynamics.
@@
Here, we focus on the post-breakup dynamics and report a novel scenario in the singular dynamics as a result of a viscosity-gravity balance: the incomplete universality where memory is partially retained. The key is a confinement, which involves a few length scales and breaks the axisymmetry of the previous cases, which allow only one length that sets the boundary conditions, i.e., a tube radius. In the present case, however, we have three independent length scales for the boundary conditions. As a result, we found the master curve, by which the self-similar dynamics is represented, does depend on the smallest length but not on the other scales, which is natural from the viewpoint of separation of scales. We show this new class of the singular dynamics by experiments with a high quality. The novel scenario of the incomplete universality on the memory in singular dynamics revealed in the present study opens a new avenue for understanding singular dynamics widely observed in nature and underscores the importance of exploration of the symmetry in singular dynamics.
@@
Figure 1: (a) Experimental setup. A metal disk of thickness $D_{0}$ ($=2.0-3.5$ mm) and radius $R$ (101010~{}10to $15$ mm) falls in the cell of thickness $D$ ($3$ to $4.5$ mm) filled with a viscous liquid of kinematic viscosity $\nu$ (1 to 50 St). The disk entrains air into the liquid, which finally detaches from the disk. The difference between $D$ and $D_{0}$ defines the liquid film thickness $e$. (b) Snapshots at breakup and after breakup illustrating the setting of axes for $(e,D_{0},R,\nu)=(0.5,3,10,1)$ in mm or St. $z=z_{c}$ is set to the origin of the $z$ coordinate.
Figure 2: Snapshots of entrainment of air by a disk into liquid, leading to breakup of a sheet of air for $(R,D_{0},e,\nu)=(10,3,0.5,10)$ in mm or St. In the top panel, overall time development is shown. In the bottom magnified snapshots, we recognize that the sharp tip at short times become rounded with time, with the thick vertical line in the middle indicates the border. The time label 0 ms corresponds to $t=t_{c}$ defined in the text.
Figure 3: (a) $z_{m}$ vs. $t^{\prime}=t-t_{c}$, where $t_{c}$ is the critical time defined in the text for various parameters $(e,D_{0},R,\nu)$ in mm or St. The data marked with a star (*) in the legend are those obtained for a different $\Delta\rho$ by using a brass disk instead of a disk of stainless steel. (b) Distinct data collapse by Eq. (1). All the data in (a) are plotted on rescaled axes, based on Eq. (1), demonstrating a clear data collapse with a quasi scaling-crossover between the regimes reasonably well characterized by the exponent $\Delta^{\prime}\simeq 0.8$ and 1/2.
Figure 4: (a1) to (a3) Temporal change of the interface $h(z,t)$ at $e=0.5$ mm for the parameter sets $(D_{0},R,\nu)=(3,10.0,5)$, $(1.0,10.0,10)$ and $(2.0,12.5,10)$ in mm or St. The error bars in capturing the interface are less than the size of markers. (b1) to (b3) The space-time dependent collapse of the shape by Eq. (2). The collapse near the tip persists for long time, while that away from the tip is observed only for short time.
Figure 5: (a) The master curves universal (and highly reproducible) for change in $D_{0}$, $R$, and $\nu=\eta/\rho$ at $e=0.5$ mm (but non-universal for change in $e$ and $\Delta\rho$, as shown below in Fig. 6 (a3)). Shapes at earlier times in Fig. 4 are compared. The data marked as (N2) are basically obtained on a different day. (b) Right branches in (a) are shown with fitting based on Eq. (3) with the axes interchanged. (c) Plot in (b) on log-log scales.
Figure 6: (a1) and (a2) Temporal change of the interface $h(z,t)$ at $\Delta\rho$ and $e$ different from those in Fig. 4 or Fig. 5: $(e,D_{0},R,\nu)=(0.5,3,10,10)$ in mm or St using a brass disk and $(1.0,2.0,10,10)$ using a stainless disk. The star (*) in (a1) [as well as in (b1), (a3) and (b3) below] indicates the data obtained for different $\Delta\rho$ as in Fig. 3. (b1) to (b2) The space-time dependent collapse of the shape by Eq. (2). (a3) The master curves non-universal for change in $e$ and $\Delta\rho$. Shapes in (b1) and (b2) at an earlier time are compared with the master curve (precisely, the fitting line) in Fig. 5 (b). (b3) $z_{m}$ vs. $t^{\prime}=t-t_{c}$. We extract from Fig. 3 (b) the data obtained for two different $\Delta\rho$’s (using the stainless and brass disks) at $(e,D_{0},R,\nu)=(0.5,3,10,10)$ in mm or St with the fitting lines in Fig. 3 (b), to clearly show the $\Delta\rho$ dependence.
I Results
I.1 Experimental
@@
In Fig. 1 (a), we show our experimental setup with explanation. The cell width and height are much larger than the length scales $R,D_{0}$, and $D$ (typically 9 and 12 cm, respectively). We use polydimethylsiloxane (PDMS) for viscous liquid of kinematic viscosity $\nu=\eta/\rho$. The density $\rho$ and the surface tension $\gamma$ are slightly depending on viscosity $\eta$ ($\rho\simeq$ $0.97$ g/cm3 and $\gamma\simeq 20$ mN/m). The density $\rho_{s}$ of the metal disk is either is 7.7 g/cm3 (stainless: SUS430) or 8.7 g/cm3 (brass) with the density difference $\Delta\rho=\rho_{s}-\rho$. The cell is fabricated with acrylic plates of thickness 5 mm, using acrylic spacers whose thickness defines the cell thickness $D$. For further details, see Appendix A.
I.2 Temporal change of the dynamics
@@
In Fig. 2, we show typical snapshots before and after breakup in the present parameter range. A thin air film is formed slightly before breakup, as confirmed by the side-view snapshots obtained for a similar parameter set in our previous study on the pre-breakup dynamics [14]. In the post breakup of the present focus, the tip is sharp when seen from the front as in the snapshots at short times (just after the breakup), but becomes rounded with time, while our focus below is mainly on the regime after the breakup but before the tip becomes clearly rounded.
I.3 Shape of the air-liquid interface $h(z,t)$
@@
We explain the definition of the shape and the setting of space-time coordinate (see Fig. 1 (b)). The shape of air-liquid interfaces formed by air entrained by the disk are seen as the dark line with a finite thickness. In the present study, the inner edge of the right (or left) line is described by: $x=h(z,t)$ (or $-h(z,t)$). For further details, see Appendix B.
@@
The space-time coordinate is determined as follows (see Fig. 1 (b)). Before breakup the function $h(z,t)$ possesses a minimum with respect to $z$, which we call ”the constriction point,” at which $(x,z)=(h_{m}(t),z_{m}(t))$, i.e., $h_{m}(t)=h(z_{m}(t),t)$. At $t=t_{c}$, topology changes: The fluid breaks into two chunks at the constriction point $(h_{m}(t_{c}),z_{m}(t_{c}))=(0,z_{c})$, where $z=z_{c}$ will be set to the origin of the $z$ coordinate. After $t=t_{c}$, the constriction point thus disappears and the dynamics of the interface of our focus is characterized by the tip point $(0,z_{m}-z_{c})=(0,z_{m})$. Experimentally, our time label 00 ms could deviate from the true $t=t_{c}$ at most 1 ms, which is set by frames per second in capturing images. See further details for Appendix B.
@@
The way of the selection the edge and the determination of time label 00 ms, detailed in Appendix B is technically important for data collapse presented below. Slight changes could deteriorate the quality of the collapse.
I.4 Dynamics of characteristic length scale
@@
In Fig. 3 (a), we present the relation between $z_{m}$ and $t$ for various conditions, in which we can confirm an excellent reproducibility of the present measurement. For example, if we closely examine the data shown by blue diamond, we see several overlapping data points, which are generally obtained on different days.
@@
with $\tau=\eta/\Delta\rho gR$, where a time label $t^{\prime}$ (which is positive at times after $t=t_{c}$) is defined as $t^{\prime}=t-t_{c}$. The data for the brass disk in Fig. 3 (b) are separately shown in Fig. 6 (b3) below, together with the corresponding data obtained by the stainless disk and the fitting lines, to show clear dependence on $\Delta\rho$. We can further confirm in Fig. 3 (b) that the slope $\Delta^{\prime}$ of the $z_{m}$-$t^{\prime}$ relation on log-log scales seem to exhibit a crossover from Regime I to Regime II: $\widetilde{z}_{m}=\widetilde{t}^{\Delta^{\prime}}$with $\widetilde{z}_{m}=z_{m}/D$ and $\widetilde{t}\simeq t^{\prime}/\tau$, where $\Delta^{\prime}\simeq 0.8$ in Regime I and $\Delta^{\prime}\simeq 1/2$ in Regime II (results of fitting are shown in the plot), although the range of regimes are limited (especially in Regime II).
I.5 Self-similarity in the interface shape dynamics
In Fig. 4 (a1) to (a3), we show temporal changes of the interfacial shape after breakup for $e=0.5$ mm but with different $\eta$, $D_{0}$, and $R$. As seen in Fig. 4 (b1) to (b3), interface shapes after rescaling both axes by $h(z,t)R/z_{m}^{2}$ and $(z-z_{m})/z_{m}$ are clearly collapsed onto a master curve, especially near the tip where $(z-z_{m})/z_{m}\,\lesssim 1$. The collapse shown in (b) implies the following scaling form, which are universal even if $\eta$, $D_{0}$, and $R$ are changed in a limited range:
@@ I.5 Self-similarity in the interface shape dynamics
As shown in Fig. 5 (a), the master curve near $t=t_{c}$ is universal (and reproducible) beyond the linear region near the tip even if parameters $D_{0}$, $R$, and $\eta$ are changed in a limited range with $e$ fixed to 0.5 mm. The universal master curve can be well described by the following expression
$\Gamma(X)=aX+bX^{2}\text{,}$ (3)
as demonstrated in Fig. 5 (b), where we obtain $a=1.26\pm 0.03$ and $b=0.65\pm 0.01$ by fitting the data at 60 ms in the range of $X$ from 0 to 5.5. The linearity of the master curve near the tip is well confirmed in Fig. 5 (c). Note that the confirmation of the collapse by Eq. (2) becomes very difficult for the shape very close to $t=t_{c}$, such as those at $t$ with $t-t_{c}\lesssim$ 10 ms, which is the order of the time resolution of our experiment limited by frames per second in capturing images. This is because $z_{m}$ becomes very small and approaches the spacial resolution of our experiment, meaning that the error in estimating $z_{m}$ becomes too large.
@@
$\Gamma(X)\simeq\left\{\begin{array}[c]{ccc}X&(X\ll 1)&\text{ for long time after }t=t_{c}\\ X^{2}&(X\gg 1)&\text{ only for short time after }t_{c}\end{array}\right.$ (4)
@@
The space-time dependent behavior of the master curve in Eq. (4) is preserved even if we change $e$ and $\Delta\rho$ in a certain range, as shown in Fig. 6 (a1) to (b2). However, the coefficients $a$ and $b$ in Eq. (3) depend on $e$ and $\Delta\rho$ as shown in Fig. 6 (a3), while they are universal for different $D_{0}$, $R$, and $\eta$ as shown in Fig. 5. In Fig. 6 (a3), the difference in the curve formed by the symbol $\triangleleft$ and the solid curve demonstrates the $\Delta\rho$ dependence, comparing the cases using brass and stainless disks with $e$ fixed to 0.5 mm. In contrast, the difference in the curve shaped by the symbol $\triangleright$ and the solid curve shows the $e$ dependence, comparing the cases of $e=1.0$ and 0.5 mm with $\Delta\rho$ fixed. We have also confirmed that the master curve at $e=0.75$ mm, for example, collapses onto neither the master curve at $e=0.5$ mm nor that at $e=1.0$ mm for a fixed $\Delta\rho$.
II Discussion
II.1 The incomplete university: partial persistence of memory

### v9

@@ Abstract
When a water drop falls from a faucet, the drop is created with the formation of an axisymmetric constriction region, which thins down to breakup. Such formation of a fluid drop has been extensively studied as a representative of the singular dynamics widely observed in nature. The singular dynamics is often self-similar, i.e., shapes at different times collapsing onto a master curve after rescaling, and the self-similar dynamics has been categorized as either universal or non-universal: the master curve is independent of or dependent on the length scales that set the initial boundary conditions, as if memory is erased or retained. Here, we focus on the post-breakup dynamics and confine the system to break the axisymmetry, introducing three length scales, which leads to a third category of incomplete universality, where memory is partially retained: the master curve could be dependent on the smallest scale but independent of the other two scales. Affecting of only the smallest length scale on the master curve underscores the importance of scale separation for the emergence of universality. The present study suggests a promising direction for the study on the singular dynamics by exploring the symmetry.
Singular dynamics observed widely in nature from the gravitational collapse of a star [1] to the formation of a drop in dripping faucets [2] have attracted scientist in many fields such as cosmology [3], hydrodynamics, physics and mathematics [4, 5, 6]. One important feature of the singular dynamics governed by partial differential equations (PDEs) is self-similarity, reported in various phenomena including fluid-jet formation [7, 8] and drop coalescence [9, 10, 11]. The self-similarity is expressed as $h(t,x)=h_{0}(t)\Gamma(x/x_{0}(t))$ for the solution $h(t,x)$ of a PDE: although the shape defined by the plot $x$ vs $h(t,x)$ for a given $t$ changes with $t$, the rescaled shape defined by the plot $X=x/x_{0}(t)$ vs $Y=h(t,x)/h_{0}(t)$ at different times collapse onto a master curve $Y=\Gamma(X)$.
The self-similar dynamics is said to be universal or losing memory, when the master curve $\Gamma(X)$ is independent of the length scales that set the initial boundary conditions (BCs). This scenario is expected from the viewpoint of separation of scales: if the physics at small scales in length and time becomes important near the singularity, the dynamics may become independent from larger scales setting the BCs, as if memory (of the initial condition) is lost. The opposite case, in which $\Gamma(X)$ is dependent on the length scales, is said to be non-universal or retaining memory. This scenario is expected when the governing equation cannot define internal length scales.
One example of the universal self-similar dynamics losing memory is the breakup of a drop surrounded by less viscous fluid, where capillarity, inertia, and viscosity all come into play [12]: While the rescaled shape at different times collapse onto a master curve, the master curve is found to be independent of the single length scale that sets the BCs, the radius of a tube to create the drop. However, the breakup of bubble surrounded by more viscous fluid was shown to be a non-universal self-similar dynamics retaining memory, where the dynamics governed by the Stokes equation describing the viscous-capillary balance [13]: The master curve does depend on the tube radius, reflecting that the Stokes equation cannot define length scales. Recently, this non-universal self-similar dynamics is found to exhibit crossover to a new universal self-similar dynamics governed by a wetting rim dynamics based on another viscous-capillary balance [14].
@@
Here, we focus on the post-breakup dynamics and report a novel scenario in the singular dynamics: the incomplete universality where memory is partially retained. The key is a confinement, which breaks the axisymmetry of the previous cases. As a result, three independent length scales are introduced for the BCs. This is in contrast with the previous axisymmetric cases, which allow only one length that sets the BCs, the tube radius. The existence of three scales leads to a novel self-similar dynamics, in which the master curve depends on the smallest length but not on the other scales: the universality is incomplete with memory partially retained. This is also expected from separation of scales if the length scale on which the collapse to a master curve is observed is comparable to the smallest length scale setting BCs but much smaller than the other scales.
We show an example of this new class of the singular dynamics by experiments with a high quality. The novel scenario of the incomplete universality on the memory in singular dynamics revealed in the present study opens a new avenue for understanding singular dynamics widely observed in nature and underscores the importance of exploration of the symmetry in singular dynamics.
@@
Figure 1: (a) Experimental setup. A metal disk of thickness $D_{0}$ (=2.03.5=2.0-3.5 mm) and radius $R$ (1010~to $15$ mm) falls in the cell of thickness $D$ (33 to $4.5$ mm) filled with a viscous liquid of kinematic viscosity $\nu$ (1 to 50 St). The disk entrains air into the liquid, which finally detaches from the disk. The difference between $D$ and $D_{0}$ defines the liquid film thickness $e$, indicated in the side view from the right edge of the front view. (b) Snapshots at breakup and after breakup illustrating the setting of axes for $(e,D_{0},R,\nu)=(0.5,3,10,1)$ in mm or St. $z=z_{c}$ is set to the origin of the $z$ coordinate.
@@
Experimental— The setup and experimental procedure are shown in Fig. 1 (a). We regard the phenomenon as a confined bubble breakup because, after air breaks up, the lower part detaches from the disk and becomes a bubble. The cell width and height (typically 9 and 12 cm, respectively) are much larger than the other length scales: the depth (or thickness) of the cell $D$, the thickness $D_{0}$ and the radius $R$ of the disk. Note that $D$ does not represent a diameter (but the cell depth), and $D_{0}$ is comparable to and a bit smaller than $D$ with $e$ $(D-D_{0})/2$ $D_{0}$. We use polydimethylsiloxane (PDMS) for viscous liquid of kinematic viscosity $\nu=\eta/\rho$. The density $\rho$ and the surface tension $\gamma$ are slightly depending on viscosity $\eta$ ($\rho\simeq$ $0.97$ g/cm3 and $\gamma\simeq 20$ mN/m). The density $\rho_{s}$ of the metal disk is either is 7.7 g/cm3 (stainless: SUS430) or 8.7 g/cm3 (brass) with the density difference $\Delta\rho=\rho_{s}-\rho$. The cell is fabricated with acrylic plates of thickness 5 mm, using acrylic spacers whose thickness defines $D$. To capture the dynamics, we analyzed the images with Image J and self-made Python codes, which were obtained with a high-speed camera (FASTCAM Mini UX 100, Photron) with a lens (Micro NIKKOR 60 mm f2.8G ED, Nikon) at 1000 to 2000 frames per second (fps). The spacial resolution of the image is typically 50 pixels per mm or more.
We here summarize important points to realize high reproducibility. We need to achieve (1) the thickness $e$ of the lubrication films on both surfaces of the disk to be equal, i.e., $e=(D-D_{0})/2$, as indicated in Fig. 1 (a), (2) the initial velocity of the metal disk to be zero at the entry (initial velocity affects the results, especially for the pre-breakup dynamics), and (3) the contact angle of the surface of the disk to be zero (the contact angle affects the results). Accordingly, (1) we set a gate with the gap $D_{0}$ at the top of cell by placing two small plates of thickness $e$ on the inside surfaces of the cell, (2) we fall the disk so that the bottom of the disk is initially in contact with the air-liquid interface, and (3) we wipe the oil on the surface of the disk after dipping to coat the surface by a thin layer of the oil.
Figure 2: Snapshots of entrainment of air by a disk into liquid, leading to breakup of a sheet of air for $(R,D_{0},e,\nu)=(10,3,0.5,10)$ in mm or St. In the top panel, overall time development is shown. In the bottom magnified snapshots, the sharp tip at short times become rounded with time, with the thick vertical line in the middle indicates the border. The time label 0 ms corresponds to $t=t_{c}$ defined in the text.
@@
Temporal change of the dynamics— In Fig. 2, we show typical snapshots before and after breakup in the present parameter range. A thin air film is formed slightly before breakup, as confirmed by the side-view snapshots obtained for a similar parameter set in our previous study on the pre-breakup dynamics [15]. In the post breakup of the present focus, the tip is sharp when seen from the front as in the snapshots at short times (just after the breakup), but becomes rounded with time, while our focus below is mainly on the regime after the breakup but before the tip becomes clearly rounded.
@@
Shape of the air-liquid interface $h(z,t)$ Figure 1 (b) explains the definition of the shape and the setting of space-time coordinate. The shape of air-liquid interfaces formed by air entrained by the disk are seen as the dark line with a finite thickness. In the present study, the inner edge of the right (or left) line is described by: $x=h(z,t)$ (or h(z,t)-h(z,t)). For further details, see Appendix A1.
@@
The space-time coordinate is determined as follows (see Fig. 1 (b)). Before breakup the function $h(z,t)$ possesses a minimum with respect to $z$, which we call ”the constriction point,” at which $(x,z)=(h_{m}(t),z_{m}(t))$, i.e., $h_{m}(t)=h(z_{m}(t),t)$. At $t=t_{c}$, topology changes: The fluid breaks into two chunks at the constriction point $(h_{m}(t_{c}),z_{m}(t_{c}))=(0,z_{c})$, where $z=z_{c}$ will be set to the origin of the $z$ coordinate. After $t=t_{c}$, the constriction point thus disappears and the dynamics of the interface of our focus is characterized by the tip point $(0,z_{m}-z_{c})=(0,z_{m})$. Experimentally, our time label 0 ms could deviate from the true $t=t_{c}$ at most 1 ms, which is set by frames per second in capturing images. See further details for Appendix A2.
@@
The way of selecting the edge and the determination of time label 0 ms, detailed in Appendix A1 and A2, are technically important for data collapse presented below. Slight changes could deteriorate the quality of the collapse.
Figure 3: (a) $z_{m}$ vs. $t^{\prime}=t-t_{c}$, where $t_{c}$ is the critical time defined in the text for various parameters $(e,D_{0},R,\nu)$ in mm or St. The data marked with a star (*) in the legend are those obtained for a different $\Delta\rho$ by using a brass disk instead of a disk of stainless steel. (b) Distinct data collapse by Eq. (1). All the data in (a) are plotted on rescaled axes, based on Eq. (1), demonstrating a clear data collapse with a quasi scaling-crossover between the regimes reasonably well characterized by the exponent $\Delta^{\prime}\simeq 0.8$ and 1/2, as indicated by the results of fitting where $y=z_{m}(t)/D$ and $x=\Delta\rho gRt^{\prime}/\eta$.
@@
Dynamics of characteristic length scale— In Fig. 3 (a), we present the relation between $z_{m}$ and $t$ for various conditions, in which we can confirm an excellent reproducibility of the present measurement. For example, if we closely examine the data shown by blue diamond, we see several overlapping data points, which are generally obtained on different days.
@@
with $\tau=\eta/\Delta\rho gR$, where a time label $t^{\prime}$ (which is positive at times after $t=t_{c}$) is defined as $t^{\prime}=t-t_{c}$ (The dependence on $\Delta\rho$ is more explicitly shown in Fig. 6 (b3) below). We can further confirm in Fig. 3 (b) that the slope $\Delta^{\prime}$ of the $z_{m}$-$t^{\prime}$ relation on log-log scales seem to exhibit a crossover from Regime I to Regime II: $\widetilde{z}_{m}=\widetilde{t}^{\Delta^{\prime}}$with $\widetilde{z}_{m}=z_{m}/D$ and $\widetilde{t}\simeq t^{\prime}/\tau$, where $\Delta^{\prime}\simeq 0.8$ in Regime I and $\Delta^{\prime}\simeq 1/2$ in Regime II (results of fitting are shown in the plot), although the range of regimes are limited (especially in Regime II).
Figure 4: (a1) to (a3) Temporal change of the interface $h(z,t)$ at $e=0.5$ mm for the parameter sets $(D_{0},R,\nu)=(3,10.0,5)$, $(1.0,10.0,10)$ and $(2.0,12.5,10)$ in mm or St. (b1) to (b3) The space-time dependent collapse of the shape by Eq. (2). The collapse near the tip persists for long time, while that away from the tip is observed only for short time.
Self-similarity in the interface shape dynamics— In Fig. 4 (a1) to (a3), we show temporal changes of the interfacial shape after breakup for $e=0.5$ mm but with different $\eta$, $D_{0}$, and $R$. As seen in Fig. 4 (b1) to (b3), interface shapes after rescaling both axes by $h(z,t)R/z_{m}^{2}$ and $(z-z_{m})/z_{m}$ are clearly collapsed onto a master curve, especially near the tip where $(z-z_{m})/z_{m}\,\lesssim 1$. The collapse shown in (b) implies the following scaling form:
@@
Figure 5: (a) The master curves universal (and highly reproducible) for change in $D_{0}$, $R$, and $\nu=\eta/\rho$ at $e=0.5$ mm (but non-universal for change in $e$ and $\Delta\rho$, as shown below in Fig. 6 (a3)). Shapes at earlier times in Fig. 4 are compared. The data marked as (N2) are basically obtained on a different day. (b) Right branches in (a) are shown with the axes interchanged, together with the results of fitting, where $y=h(z,t)R/z_{m}^{2}$ and $x=(z-z_{m})/z_{m}$.
As shown in Fig. 5 (a), the master curve near $t=t_{c}$ collapses well (and thus is reproducible) beyond the linear region near the tip up roughly to $(z-z_{m})/z_{m}=3$ even if parameters $D_{0}$, $R$, and $\eta$ are changed in a limited range with $e$ fixed to 0.5 mm. The universal master curve seems to be linear near the tip but quadratic away from the tip. In fact, we can show the master curve is well described by $\Gamma(x)=ax+bx^{2}$ with $a=1.8\pm 0.2$ and $b=0.4\pm 0.05$ as shown in Fig. 5 (b) by a two-step fitting: We first determine $a$ by fitting to the linear region near the tip (on log-log scales) and then determine $b$ by fitting to the quadratic region after subtracting the linear component. Note that the confirmation of the collapse by Eq. (2) becomes very difficult for the shape very close to $t=t_{c}$, such as those at $t$ with $t-t_{c}\lesssim$ 10 ms, which is the order of the time resolution of our experiment limited by frames per second in capturing images. This is because $z_{m}$ becomes very small and approaches the spacial resolution of our experiment, meaning that the error in estimating $z_{m}$ becomes too large.
@@
[…]
In other words, in Figs. 4 (and 6 below), the master curve is confirmed in the spacial region where the rescaled curves at the earliest time and that at the second earliest collapse: The spacial region of the master curve confirmed in this way is up roughly to $(z-z_{m})/z_{m}=3$.
Figure 6: (a1) and (a2) Temporal change of the interface $h(z,t)$ at $\Delta\rho$ and $e$ different from those in Fig. 4 or Fig. 5: $(e,D_{0},R,\nu)=(0.5,3,10,10)$ in mm or St using a brass disk and $(1.0,2.0,10,10)$ using a stainless disk. The star (*) in (a1) [as well as in (b1), (a3) and (b3) below] indicates the data obtained for different $\Delta\rho$ as in Fig. 3. (b1) to (b2) The space-time dependent collapse of the shape by Eq. (2). (a3) The master curves non-universal for change in $e$ and $\Delta\rho$. Shapes in (b1) and (b2) are compared with a shape in Fig. 4 (b2) at an earlier time. (b3) $z_{m}$ vs. $t^{\prime}=t-t_{c}$. We extract from Fig. 3 (b) the data obtained for two different $\Delta\rho$’s (using the stainless and brass disks) at $(e,D_{0},R,\nu)=(0.5,3,10,10)$ in mm or St with the fitting lines in Fig. 3 (b), to clearly show the $\Delta\rho$ dependence. The results of the fitting are also shown with $y=z_{m}(t)/D$ and $x=\Delta\rho gRt^{\prime}/\eta$.
@@
The space-time dependent behavior of the master curve in Eq. (3) is preserved even if we change $e$ and $\Delta\rho$ in a certain range, as shown in Fig. 6 (a1) to (b2). However, the master curve defined by shapes at earlier times in the range up roughly to $(z-z_{m})/z_{m}=3$ clearly depends on $e$ and $\Delta\rho$ as demonstrated in Fig. 6 (a3), where we compare the shape at one early time with those obtained for a different $e$ and a different $\Delta\rho$. We have also confirmed that the master curve at $e=0.75$ mm, for example, collapses onto neither the master curve at $e=0.5$ mm nor that at $e=1.0$ mm for a fixed $\Delta\rho$.